# Investigating the Electroweak Phase Transition with a Real Scalar Singlet at a Muon Collider

### v1

@@ 4.7 $2q2\nu_{l}$ Final states
Unlike the $2l2\nu_{l}$ case, this final state can originate from the $S$ decay into $Z$-bosons, $S\rightarrow ZZ,$ together with all possible backgrounds resulting from $\mu^{+}\mu^{-}\rightarrow\mu^{+}\mu^{-}(\nu_{\mu}\overline{\nu}_{\mu})2q2\nu_{l}$. Similar to the previous channel, it would be impossible to directly compute the invariant mass owing to the missing energy carried away by the neutrinos. However, we can employ the same formula as Eq. (42) to evaluate it, with a slight replacement of $l^{+}l^{-}\rightarrow nJ$, where all the jets contained in the event are sourced from a single $Z$ boson, and $\Delta$ becomes a function of the forward muons ($q_{1},q_{2}$) and the all the jets ($p_{1},p_{2},\cdots p_{n}$).
@@
In order to accept an event, we further require that it must contain at least one jet pair, $J\overline{J}$, with $0.5M_{Z}\leq M(J\overline{J})\leq 1.4M_{Z})$ for $P_{T}(J,\overline{J})\geq\frac{1}{5}M_{Z}$. We then applied cuts on $P_{T}(J\overline{J}),\eta(J\overline{J})$, obtained through a direct signal-background comparison, as shown in Fig. 19 (A,B) for $M_{s}=550$ GeV. These secondary cuts were automated for each $M_{s}$ value, as illustrated in Table 8 for a sample of selected $M_{s}$ values. The $P_{T}$ bins were selected by requiring $N_{s}\geq 1.2N_{b}$ and $N_{s}\geq 1.4N_{b}$ for the $\eta$ bins. Similar to all previous channels, the uncertainty in $\varepsilon_{s},\varepsilon_{b}$ was computed using Eq.(36) by repeating the preceding analysis for a different invariant mass range, $0.85M_{X}\leq M(l^{+}l^{-}+\Delta)\leq 1.6M_{X}$.
@@
This analysis scheme greatly suppressed the background, yielding a background efficiency of $\varepsilon_{b}\sim 1\%$ on average, whereas the signal efficiency $\varepsilon_{s}$ exceeded $25\%$ on average. The corresponding sensitivity plot for this final state is shown in Fig. 20 for a luminosity of $L=10^{4}~{}\mathrm{fb}^{-1}$.
@@ 4.8 $l\nu_{l}2q$ Final states
This final state is another gauge boson-initiated channel that only arise from $S\longrightarrow W^{+}W^{-}\rightarrow l^{+}\nu_{l}(l^{-}\tilde{\nu_{l}})q\overline{q}$. All backgrounds contributing to this channel that could be generated at the muon collider from $\mu^{+}\mu^{-}\longrightarrow mu^{+}\mu^{-}l^{+}\nu_{l}(l^{-}\tilde{\nu_{l}})q\overline{q}$ were considered. The invariant mass of the fully visible final states can also be evaluated using Eq. 42, but now ‘$\Delta$’ will be a function of an extra lepton vector (i.e. $p_{3}$) which results from $W$-boson decay,
@@
$M(l^{\pm}+nJ)$ runs over all possible combinations of a single charged lepton in the final state, with $n$ jets available in the event. This includes forward high-energy scattered muons, so a secondary peak around $M_{S}\approx\sqrt{s}$ should be expected, and could be ignored for the reasons discussed earlier. In fact, this will not even be seen in this case due to the jets largely outnumbering the the forward muons, and hence the secondary peak at $M_{S}=\sqrt{s}$ will be smeared away, which would be further enhanced by only considering charged leptons with $P_{T}(l^{\pm})\geq\frac{1}{5}M_{W}$ that shall suppress the forward muons contributions. This approach generated well-defined invariant mass distributions for all the examined $M_{s}$ values. as can be seen in Fig. 21 (C), which suggests narrower invariant mass range compared to the previous channel, $0.85M_{s}\leq M(l^{\pm}+nJ+\Delta)\leq 1.5M_{s}$. If the event passes this primary cut, we then check whether the jet invariant mass peaks around $W$-boson mass, $0.75M_{W}\leq M(nJ)\leq 1.25M_{W}$. If it does, we then construct all the possible jet pairs with $P_{T}\geq\frac{1}{5}M_{W}$. At least one jet pair, $J\overline{J}$, is required in order to pass this event to the secondary analysis. In this step we checked for the observables $P_{T}(J\overline{J}),\eta(J\overline{J})$, the regions where the signal outnumbers the background, keeping only the bins where $N_{s}\geq 1.1N_{b}$ for $P_{T}(J\overline{J})$ and $N_{s}\geq 1.2N_{b}$ for the $\eta(J\overline{J})$. This scanning procedure is automated for each $M_{s}$ value as shown in Fig. 21 (A,B) for $M_{s}=600$ GeV, and for other different scalar mass values in Table.9. We finally account for the fluctuations in the measured efficiencies by repeating the previous analysis for different invariant mass range, $0.8M_{s}\leq M(l^{\pm}+nJ+\Delta)\leq 1.65M_{s}$, and then compute $U(\sigma_{s})$ using Eq. (36).
@@
Our analysis successfully suppressed backgrounds in comparison to signal events, with $\varepsilon_{b}\sim 10:40\%$, and $\varepsilon_{s}\sim 25:88\%$. This channel shows a sensitivity in $\varepsilon_{s},\varepsilon_{b}$ to the scalar mass value, as both tend to grow for higher $M_{s}$ values, a feature that was not observed in the other channels. This feature is reflected in the sensitivity plot in Fig. 22, where the uncertainty in the signal cross-section converges towards higher values of the scalar mass as $\varepsilon_{s}$ becomes larger than $50\%$, which guarantees that the uncertainties will be smaller.
4.9 Exclusion Plots
Utilizing the aforementioned phenomenological analyses, we explored the parameter-space points that fulfill the SFOEWPT conditions using the derived sensitivity plots, as illustrated in Fig. 23.
@@ 4.9 Exclusion Plots
It is evident that channels containing quarks in the final state can explore a larger portion of the xSM SFOEWPT parameter space. This is primarily attributed to the higher branching ratios of the Higgs bosons and vector bosons to the jets. The relative absence of substantial QCD backgrounds at a muon collider, in comparison to hadron colliders, enables the reconstruction of jets via a comparatively straightforward approach. This factor facilitates efficient background suppression, even for channels with invisible final states. This same rationale supports the pursuit of potential precision tests of the Higgs boson’s self-couplings at a muon collider, and the possible confirmation or exclusion of the SFOEWPT scenario. We reserve such an investigation for future studies, along with the examination of potential kinematical differences between jets originating from $S\rightarrow hh$ and $S\rightarrow ZZ(WW)$, which may be related to the $\mathbb{Z}_{2}$-nature of xSM extended potential.
@@
Figure 24: (A) shows the combined $WW,\,ZZ,\,hh\rightarrow nJ$ final states at a muon collider with $\sqrt{s}=3$ TeV and $L=10^{5}$ ab-1. (B) shows the FCC-hh exclusion curves for the same channels at a much higher energy $\sqrt{s}=100$, TeV and much smaller luminosity $L=20\,~{}\mathrm{fb}^{-1}$.
Figure 23 presents the currently viable parameter-space points that satisfy the SFOEWPT conditions for both the liberal and conservative categories discussed in Section 2 in conjunction with the sensitivity plots obtained from the previous subsections. Figure 23 (A) summarizes all the different final states that could be obtained from $WW$ decay, and Fig. 23 (B,C) for $ZZ$ and $hh$, respectively, at a luminosity of $10^{5}~{}\mathrm{fb}^{-1}$. [^7] The exclusion curve is proportional to $\frac{1}{\sqrt{L}}$ (see Eq. (35)), so we have chosen an order of magnitude higher luminosity than the value used in the previous exclusion curves in order to explore a larger parameter space volume. The preferred channels, that is, jet-rich channels, have consequently been isolated in Fig. 24, as they represent the primary limiting channels in a muon collider. In comparison to the FCC-hh potential for searches, depicted in Fig. 24 (B), the muon collider appears to be capable of excluding the majority of points that could be examined at the FCC-hh at significantly lower center-of-mass energies, particularly for $S\rightarrow XX\rightarrow 4q$.
@@ 5 Conclusions
In this study, we investigated the potential of a high-energy muon collider to explore electroweak phase transitions in the context of the Standard Model extended by a real scalar singlet. Our analysis focused on examining the occurrence of a strong first-order electroweak phase transition, which is essential for explaining baryogenesis and may result in observable gravitational waves.
As the initial step of our investigation, we identified Lagrangian parameter values for which xSM can produce a strong first-order phase transition that satisfies the sphaleron critical conditions for baryogenesis. These parameter values span a broad parameter space for the xSM. Subsequently, assuming a future muon collider with center-of-mass energy $\sqrt{s}=3$ TeV, we assessed the sensitivity to a scalar singlet through various final states. Our findings indicate that the muon collider can be an effective tool for probing the xSM parameter space if high luminosities can be approached, where the clean environment of a muon collider allows for precise measurements, particularly producing a scalar singlet via vector boson fusion (VBF). We primarily examined the direct production of the scalar singlet and its decay into all possible final states, which offers a promising avenue for identifying new scalar states. We found that decay channels rich with jets in the final states are the most promising for excluding FOEWPT parameter space points at the muon collider. Indirectly, deviations in the Higgs boson couplings from the SM expectations could provide further evidence for the presence of a scalar singlet, which would be valuable to explore further in the muon collider. With precision measurements, the zero-temperature part of the effective potential can be reconstructed at the muon collider, whereas gravitational wave observations can provide additional information on thermal dynamics. Scalar singlet-enhanced electroweak phase transition can also generate a signal in gravitational wave detectors, which enables the muon collider to provide complementary information for gravitational wave detectors.
Our findings highlight the unique advantages of a muon collider in exploring electroweak symmetry breaking and potentially new physics beyond the Standard Model. The ability to resolve the nature of the EWPT in such a collider is of paramount importance, especially for advancing our understanding of early universe dynamics and the mechanisms behind baryogenesis, providing significant advancement in our understanding of electroweak symmetry breaking and its implications in cosmology.
@@ Acknowledgements.
The research of C.B. is supported by the Australian Research Council via DP210101636, DP220100643, and LE210100015. A.P. acknowledges support by the National Science Foundation under Grant No. PHY 2210161.

### v2

@@ 4.7 $2q2\nu_{l}$ Final states
Unlike in the $2l2\nu_{l}$ case, this final state can originate from the $S$ decay into $Z$-bosons, $S\rightarrow ZZ,$ together with all possible backgrounds resulting from $\mu^{+}\mu^{-}\rightarrow\mu^{+}\mu^{-}(\nu_{\mu}\overline{\nu}_{\mu})2q2\nu_{l}$. Similarly to the previous channel, it would be impossible to directly compute the invariant mass owing to the missing energy carried away by the neutrinos. However, we can employ the same formula as Eq. (42) to evaluate it, with a slight replacement of $l^{+}l^{-}\rightarrow nJ$, where all the jets contained in the event are sourced from a single $Z$ boson, and $\Delta$ becomes a function of the forward muons ($q_{1},q_{2}$) and all the jets ($p_{1},p_{2},\cdots p_{n}$).
@@
In order to accept an event, we further require that it must contain at least one jet pair, $J\overline{J}$, with $0.5M_{Z}\leq M(J\overline{J})\leq 1.4M_{Z})$ for $P_{T}(J,\overline{J})\geq\frac{1}{5}M_{Z}$. We then applied cuts on $P_{T}(J\overline{J}),\eta(J\overline{J})$, obtained through a direct signal-background comparison, as shown in Fig. 19 (A,B) for $M_{s}=550$ GeV. These secondary cuts were automated for each $M_{s}$ value, as illustrated in Table 8 for a sample of selected $M_{s}$ values. The $P_{T}$ bins were selected by requiring $N_{s}\geq 1.2N_{b}$ and $N_{s}\geq 1.4N_{b}$ for the $\eta$ bins. Similarly to all previous channels, the uncertainty in $\varepsilon_{s},\varepsilon_{b}$ was computed using Eq.(36) by repeating the preceding analysis for a different invariant mass range, $0.85M_{X}\leq M(l^{+}l^{-}+\Delta)\leq 1.6M_{X}$.
@@
This analysis scheme greatly suppressed the background, yielding a background efficiency of $\varepsilon_{b}\sim 1\%$ on average, while the signal efficiency $\varepsilon_{s}$ exceeded $25\%$ on average. The corresponding sensitivity plot for this final state is shown in Fig. 20 for a luminosity of $L=10^{4}~{}\mathrm{fb}^{-1}$.
@@ 4.8 $l\nu_{l}2q$ Final states
This final state is another gauge boson-initiated channel that arises only from $S\longrightarrow W^{+}W^{-}\rightarrow l^{+}\nu_{l}(l^{-}\tilde{\nu_{l}})q\overline{q}$. All backgrounds contributing to this channel that could be generated at the muon collider from $\mu^{+}\mu^{-}\longrightarrow mu^{+}\mu^{-}l^{+}\nu_{l}(l^{-}\tilde{\nu_{l}})q\overline{q}$ were considered. The invariant mass of the fully visible final states can also be evaluated using Eq. 42, but now ‘$\Delta$’ will be a function of an extra lepton vector (i.e. $p_{3}$) which results from $W$-boson decay,
@@
$M(l^{\pm}+nJ)$ runs over all possible combinations of a single charged lepton in the final state, with $n$ jets available in the event. This includes forward high-energy scattered muons, so a secondary peak around $M_{S}\approx\sqrt{s}$ should be expected, and could be ignored for the reasons discussed earlier. In fact, this will not even be seen in this case because the jets largely outnumber the forward muons, and hence the secondary peak at $M_{S}=\sqrt{s}$ will be smeared away, which would be further enhanced by only considering charged leptons with $P_{T}(l^{\pm})\geq\frac{1}{5}M_{W}$ that shall suppress the forward muon contributions. This approach generated well-defined invariant mass distributions for all the examined $M_{s}$ values, as can be seen in Fig. 21 (C), which suggests a narrower invariant mass range compared to the previous channel, $0.85M_{s}\leq M(l^{\pm}+nJ+\Delta)\leq 1.5M_{s}$. If the event passes this primary cut, we then check whether the jet invariant mass peaks around the $W$-boson mass, $0.75M_{W}\leq M(nJ)\leq 1.25M_{W}$. If it does, then we construct all possible jet pairs with $P_{T}\geq\frac{1}{5}M_{W}$. At least one pair of jets, $J\overline{J}$, is required in order to pass this event to the secondary analysis. In this step, we check for the observables $P_{T}(J\overline{J}),\eta(J\overline{J})$, the regions where the signal outnumbers the background, keeping only the bins where $N_{s}\geq 1.1N_{b}$ for $P_{T}(J\overline{J})$ and $N_{s}\geq 1.2N_{b}$ for $\eta(J\overline{J})$. This scanning procedure is automated for each $M_{s}$ value as shown in Fig. 21 (A,B) for $M_{s}=600$ GeV, and for other different scalar mass values in Table9. We finally account for the fluctuations in the measured efficiencies by repeating the previous analysis for a different invariant mass range, $0.8M_{s}\leq M(l^{\pm}+nJ+\Delta)\leq 1.65M_{s}$, and then compute $U(\sigma_{s})$ using Eq. (36).
@@
Our analysis successfully suppressed backgrounds in comparison to signal events, with $\varepsilon_{b}\sim 10:40\%$, and $\varepsilon_{s}\sim 25:88\%$. This channel shows a sensitivity in $\varepsilon_{s},\varepsilon_{b}$ to the scalar mass value, as both tend to grow for higher $M_{s}$ values, a feature that was not observed in the other channels. This feature is reflected in the sensitivity plot in Fig. 22, where the uncertainty in the signal cross section converges towards higher values of the scalar mass as $\varepsilon_{s}$ becomes larger than $50\%$, which guarantees that the uncertainties will be smaller.
5 Results
Using the phenomenological analyses described above, we explored the parameter space points that fulfill the SFOEWPT conditions using the derived sensitivity plots, as illustrated in Fig. 23.
@@ 5 Results
It is evident that channels containing quarks in the final state can explore a larger portion of the xSM SFOEWPT parameter space. This is primarily attributed to the higher branching ratios of the Higgs bosons and vector bosons to the jets. The relative absence of substantial QCD backgrounds at a muon collider, in comparison to hadron colliders, enables the reconstruction of jets via a comparatively straightforward approach. This factor facilitates efficient background suppression, even for channels with invisible final states. This same rationale supports the pursuit of potential precision tests of the Higgs boson self-couplings at a muon collider, and the possible confirmation or exclusion of the SFOEWPT scenario. We reserve such an investigation for future studies, along with the examination of potential kinematical differences between jets originating from $S\rightarrow hh$ and $S\rightarrow ZZ(WW)$, which may be related to the $\mathbb{Z}_{2}$ nature of the extended potential of the xSM.
@@
Figure 24: (A) shows the combined $WW,\,ZZ,\,hh\rightarrow nJ$ final states at a muon collider with $\sqrt{s}=3$ TeV and $L=10^{5}$ ab-1. (B) shows the FCC-hh exclusion curves for the same channels at a much higher energy $\sqrt{s}=100$, TeV and much smaller luminosity $L=20\,~{}\mathrm{fb}^{-1}$. (C) represents the discovery plot for $\mathcal{S}=5$.
Figure 23 presents the currently viable parameter space points that satisfy the SFOEWPT conditions for both the liberal and conservative categories discussed in Section 2 in conjunction with the sensitivity plots obtained from the previous subsections. Figure 23 (A) summarizes all the different final states that could be obtained from $WW$ decay, and Fig. 23 (B,C) for $ZZ$ and $hh$, respectively, at a luminosity of $10^{5}~{}\mathrm{fb}^{-1}$. [^7] The exclusion curve is proportional to $\frac{1}{\sqrt{L}}$ (see Eq. (35)), so we have chosen an order of magnitude higher luminosity than the value used in the previous exclusion curves in order to explore a larger parameter space volume. The preferred channels, that is, jet-rich channels, have consequently been isolated in Fig. 24, as they represent the primary limiting channels in a muon collider. Compared to the FCC-hh potential for searches, depicted in Fig. 24 (B), the muon collider appears to be capable of excluding the majority of points that could be examined at the FCC-hh at significantly lower center-of-mass energies, particularly for $S\rightarrow XX\rightarrow 4q$.
In the event of a positive detection of the real singlet at the proposed future muon collider, it will still be possible to exclude a significant portion of FOEWPT parameter space points. This accounts for a modification of the statistical significance in Eq.(35) from $\mathcal{S}=2$ to $\mathcal{S}=5$, which results in a consistent and slight upward shift of the sensitivity curves across all channels. This result is illustrated in Fig. 24 (c). From that discovery plot, we can further estimate, in the event of a positive detection, the anticipated mixing between the two scalars. As discussed in Section 3, the potential in Eq. (3) results in mixing between the SM Higgs and the new scalar, which is a desirable characteristic for enhancing the FOEWPT. This mixing will scale all the Higgs couplings by $\cos\theta$ papaefstathiou2022electro .
Figure 25: The limits in the Higgs-real scalar mixing extracted from the different decay channels that could be probed at the muon collider.
In Ref. Han_2021 , the authors discussed the sensitivity of the Higgs boson’s properties to new physics, particularly its couplings to weak gauge bosons, and the Higgs trilinear coupling. In this investigation, the source of the ‘new physics’ is the real singlet extension, which inevitably scales all of the Higgs boson’s couplings. Of particular importance for this study is the modification of the Higgs trilinear coupling due to this extension, given its role in governing the Higgs potential and, consequently, the type and strength of the EWPT. This work is currently in progress. However, utilizing the discovery plots in Fig. 24 (C), it is possible to extrapolate certain limits on the mixing angle between the two scalars, which may be observed in current and future precision tests of the Higgs properties. Following ref. papaefstathiou2022electro , the mixing angle between the two scalars can be derived from,
[…]
where, $\sigma_{\text{\tiny BSM}}$ represents the expected cross section obtained from our analysis in Section. 4, and $\sigma_{\text{\tiny SM}}$ denotes the SM cross section for each final state topology computed using MG5_aMC, setting $m_{h}=m_{s}$ ($\sin\theta=1$). Through this analysis for the muon collider, it is anticipated that $0.02\leq\sin\theta\leq 0.15$ in the event of a positive discovery. This result can already exclude some of the valid parameter space points for FOEWPT, as each set predicts a specific value for the mixing angle, which is illustrated in Fig.25 for both the conservative and liberal categories.
@@ 6 Conclusions
In this study, we investigated the potential of a high-energy muon collider to explore electroweak phase transitions in the context of the Standard Model extended by a real scalar singlet. Our analysis focused on examining the occurrence of a strong first-order electroweak phase transition, which is essential for explaining baryogenesis, and may result in observable gravitational waves.
As the initial step of our investigation, we identified Lagrangian parameter values for which xSM can produce a strong first-order phase transition that satisfies the sphaleron critical conditions for baryogenesis. These parameter values span a broad parameter space for the xSM. Subsequently, assuming a future muon collider with center-of-mass energy $\sqrt{s}=3$ TeV, we assessed the sensitivity to a new scalar singlet-like particle through various final states. Our findings indicate that the muon collider can be an effective tool for probing the xSM parameter space if high luminosities can be approached, where the clean environment of a muon collider allows for precise measurements, particularly producing a scalar singlet-like particle via vector boson fusion (VBF). We primarily examined the direct production of the new scalar and its decay into all viable final states, which offers a promising avenue for its discovery or exclusion. We found that decay channels rich with jets in the final states are the most promising for excluding FOEWPT parameter space points at the muon collider. Indirectly, deviations in the Higgs boson couplings from the SM expectations could provide further evidence for the presence of a new scalar singlet-like particle, which would be valuable to explore further in the muon collider. With precision measurements, the zero-temperature part of the effective potential can be reconstructed at the muon collider, whereas gravitational wave observations can provide additional information on thermal dynamics. The scalar singlet-enhanced electroweak phase transition can also generate a signal in gravitational wave detectors, which enables the muon collider to provide complementary information for gravitational wave detectors.
Our findings highlight the unique advantages of a muon collider in exploring electroweak symmetry breaking and potentially new physics beyond the Standard Model. The ability to resolve the nature of the electroweak phase transition in such a collider is of paramount importance, especially for advancing our understanding of early universe dynamics and the mechanisms behind baryogenesis, providing significant advancement in our understanding of electroweak symmetry breaking and its implications in cosmology.
@@ Acknowledgements.
The research of C.B. is supported by the Australian Research Council through DP210101636, DP220100643, and LE210100015. A.P. acknowledges support by the National Science Foundation under Grant No. PHY 2210161.